OpenStudy (anonymous):

need help 5x + 2y = 10 3x + 2y = 6

OpenStudy (jdoe0001):

specific postings helps

OpenStudy (anonymous):

Solve the system algebraically

TheSmartOne (thesmartone):

Welcome to OpenStudy!

TheSmartOne (thesmartone):

Do you have any clue on how to solve your problem?

OpenStudy (anonymous):

no not really

TheSmartOne (thesmartone):

Do you know how to solve the equations by elimination?

OpenStudy (anonymous):

no

TheSmartOne (thesmartone):

So basically to solve be elimination we just subtract the 2 equations to get rid of one variable and solve for the other variable. Then we plug in the value for the variable we got into one of the equations and we get the value of the last variable.

OpenStudy (anonymous):

2x+0=4

TheSmartOne (thesmartone):

exactly!

TheSmartOne (thesmartone):

So now solve for x! What do you get 2x+0 is nothing but 2x So 2x=4 What is x=?

OpenStudy (anonymous):

x=2

TheSmartOne (thesmartone):

So if x=2 then we should just plug it in back to the original question and solve for y. 3x + 2y = 6 3(2)+2y=6 Solve for y. You could even plug in x=2 into the other equation to solve for y. :D

OpenStudy (anonymous):

y=0